[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jun 15, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jun 15 18:28:39 UTC 2016


HSU Shift Log for Wed Jun 15 2016 14:28:33 GMT-0400 (EDT)

Shift: H

Forecaster: R Pasch

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: ASCAT-B has not been available today in N-AWIPS. The  
FNMOC page showed that it missed the area where a center of EP93 might  
exist.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160615/b35d1fc4/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list